Two loudspeakers emit sound waves along the x-axis
Answer #1

Highest intensity = 15cm
Lowest intensity = 62cm

that is 62-15 = 47 = \lambda/ 2

\lambda   = 47*2 = 94 cm ..................answer

From 62 cm to another point of high intensity = 62 + 47 = 109 cm